# OPPO: Accelerating PPO-based RLHF via Pipeline Overlap

### v1

@@ Abstract
Proximal Policy Optimization (PPO)-based reinforcement learning from human feedback (RLHF) is a widely adopted paradigm for aligning large language models (LLMs) with human preferences. However, its training pipeline suffers from substantial inefficiencies due to sequential multi-model dependencies (e.g., reward model depends on actor outputs) and long-tail response lengths, where a few long responses straggle the stage completion. We present OPPO, a novel, lightweight, and model-agnostic PPO-based RLHF framework that improves training efficiency by overlapping pipeline execution. OPPO introduces two novel techniques: (1) Intra-step overlap, which streams upstream model outputs (e.g., actor model) in right-sized chunks, enabling the downstream model (e.g., reward) to begin prefill while the upstream continues decoding; and (2) Inter-step overlap, which adaptively overcommits a few prompts and defers long generations to future steps, mitigating tail latency without discarding partial work. OPPO integrates easily with existing PPO implementations with a few lines of code change. Extensive evaluations show that OPPO accelerates PPO-based RLHF training by $1.8\times$$2.8\times$ and improves GPU utilization by $1.4\times$$2.1\times$ without compromising training convergence.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human preferences. Among RLHF methods, Proximal Policy Optimization (PPO) (Schulman et al., 2017) has been the de facto standard due to its training stability and flexibility across diverse reward models and objectives. PPO underpins the training of state-of-the-art LLMs such as GPT-4 (OpenAI et al., 2024), Gemma (Team et al., 2024), and LLaMA (Touvron et al., 2023; Grattafiori et al., 2024). A standard PPO-based RLHF pipeline involves four models: an actor (policy), a critic (value function), a reference policy (for KL regularization), and a reward model trained on human-labeled preferences. Each training step consists of three sequential stages: (1) Generation: the actor generates responses to prompts; (2) Scoring: responses are evaluated by the critic, reference, and reward models; and (3) Training: the actor, critic, and policy models are updated using advantage estimates and gradients.
@@
Despite its effectiveness, PPO-based RLHF faces large training inefficiencies rooted in its multi-model dependency. Running and coordinating four LLMs imposes substantial resource requirements, and each stage is constrained by its slowest component. For example, the actor model’s generation suffers from severe long-tail latency: a few long responses can delay downstream stages, such as the reward and value models, leading to idle resources and poor training throughput of the pipeline. As LLMs grow larger and context lengths increase, these bottlenecks worsen,  (Grattafiori et al., 2024) making PPO-based RLHF increasingly costly to train (§2.2).
@@
Recent advances tackle PPO-based RLHF inefficiencies from both algorithmic and system angles. On the algorithmic side, methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Group-Relative Policy Optimization (GRPO) (Shao et al., 2024) remove components like the value or reward model. However, these approaches often suffer from instability due to sparse rewards, requiring many rollouts to capture intrinsic advantages, and face task-specific reward design challenges (Feng et al., 2024; Fisch et al., 2025; Chen et al., 2024). Asynchronous RLHF (Noukhovitch et al., 2025) reduces pipeline dependencies but introduces staleness, which can harm convergence (§2.2). In contrast, system-level approaches, such as RLHFuse (Zhong et al., 2025), AReal (Fu et al., 2025), and Verl (Sheng et al., 2025), improve throughput via fine-grained parallelism, for example by collocating models to reduce communication overhead or dynamically scaling GPU resources to match workload demands.
@@
In this paper, we explore a complementary opportunity to accelerate PPO-based RLHF: maximizing execution overlap in the training pipeline. We introduce two novel insights: (1) Intra-step Overlap, which streams tokens from the actor to downstream models, enabling generation and scoring stages to overlap the execution without altering the generated response; and (2) Inter-step overlap, which selective overcommits a few prompts per batch and selects faster completions, deferring stragglers to future iterations to hide tail latency without wasting partial generation.
@@
In this paper, we present OPPO, a lightweight PPO-based RLHF training framework that improves training efficiency via pipeline overlap, minimizing idle time without compromising convergence. OPPO novelly address the aforementioned challenges:
  • Intra-step Overlap: While the actor generates responses, OPPO streams newly generated tokens to downstream models (e.g., reward model) in adaptive chunks. This enables incremental prefilling and overlaps the generation and scoring stages. Chunk sizes are automatically adjusted online, based on the leftover resource, to balance overlap against resource contention, preserving algorithm correctness and stability (§3.1).
@@
  • Generalized and lightweight: Our evaluations show that OPPO achieves $1.8\times$$2.8\times$ speedup and improves GPU utilization by $1.4\times$$2.1\times$ for PPO-based RLHF with only a few lines of code changes, and generalizes to other paradigms such as DPO with similar benefits (§4).
@@
Figure 1: (a) In the existing paradigm, the scoring stage starts to process until that response is fully generated. In contrast, (b) the OPPO paradigm interleaves scoring with generation without altering the final responses (intra-step overlap), and carries unfinished overcommitted sequences into the next iteration (inter-step overlap). A batch size of 4 and an overcommitment degree of 2 in illustrations.
@@ 2 Background and Motivation
Figure 2: PPO-based RLHF faces (a) varying resource demands across pipeline stages, and (b) response lengths across rollouts, both of which can produce stragglers that prolong step execution. (c) Existing approaches for asynchronous training risk harming convergence.
@@ 2.1 Background: PPO-based RLHF
Figure 1(a) depicts a single step of a standard PPO-based RLHF pipeline. Given a batch of prompts, the actor model generates output sequences. These are then scored by a reward model, producing scalar rewards that reflect alignment with human preferences. A reference model, typically a frozen copy of the base pretrained model, computes a KL divergence penalty that regularizes the update, discouraging the new policy from drifting too far from the original distribution.
@@
where $r_{t}$ is the reward at step $t,V\left(s_{t}\right)$ is the estimated value of state $s_{t},\gamma$ is the discount factor, and $\lambda$ is the generalized advantage estimation (GAE) parameter. Every step actor model is updated by optimizing the clipped surrogate objective:
@@ 2.2 Training Inefficiency of PPO-based RLHF
Each model in the RLHF pipeline (e.g., actor, reward, and value models) exhibits distinct computational characteristics. Figure 2(a) compares GPU utilization across three GPUs (A40, A100, H200). Response generation in the actor model is memory-intensive due to autoregressive per-token decoding, resulting in low GPU utilization (<40%), whereas scoring and training stages are relatively compute-intensive (e.g., due to long-context prefilling in scoring). This heterogeneous resource utilization highlights how mismatched compute demands across stages create idle GPU time, motivating designs to scavenge unused resources.
@@
Each step involves updating model weights. A natural strategy to improve throughput is to tolerate training staleness, where the reward model evaluates actor outputs from previous steps instead of synchronizing in real time, as in AReal  (Fu et al., 2025). However, as shown in Figure 2(c), asynchronous training (e.g., with staleness 5) can not only slows step-to-reward convergence but also lowers the post-training model quality, emphasizing the need for careful inter-step synchronization in RLHF pipelines.
@@ 3.1 Overlapping Intra-step Training Pipeline
By dividing actor generation into chunks and streaming them to the reward model, OPPO overlaps the actor decoding stage with the reward prefilling stage, hiding latency and reducing execution bubbles. This design naturally benefits setups where models are placed on separate GPUs, but also improves efficiency when models are colocated, due to their mismatched compute demands (Figure 2(a)). To realize intra-step overlap, OPPO partitions actor outputs into right-sized chunks and streams each chunk to the reward model as it is generated. Scoring proceeds progressively within each PPO step: while the actor decodes the $k$-th chunk, the reward model concurrently processes the prefilling of (k1)(k-1)-th chunk. At the end of the step, the reward model completes the last-chunk prefilling and generates the score based on the entire sequence, whose previous chunks have already be processed.
Importantly, this streaming does not alter the response generation $y_{i}$, the policy log-probabilities, or the critic/value terms used in computing the advantage $\hat{A}(y_{i})$. Formally, letting $y_{i}$ be the full response and $y_{i}^{(1)},\dots,y_{i}^{(T_{i})}$ its prefixes with $y_{i}^{(T_{i})}=y_{i}$, the streamed gradient estimator is
@@ 3.2 Overlapping Inter-step Training Pipeline
While intra-step overlap improves efficiency within a single PPO step, it does not fully address tail latency caused by the heterogeneous response lengths of prompts in a batch. Here, the response must complete generation before its reward scoring and subsequent policy updates. Due to the long-tailed distribution of generation lengths, a few slow prompts can delay the entire step. This motivates an inter-step design that allows overlapping across PPO steps without hurting convergence.
@@
The overall procedure, combining intra- and inter-step overlap, is summarized in Algorithm 1, where the buffer holds up to $B+\Delta$ sequences, and generation proceeds in parallel with intra-step streaming. The threshold $\Delta$ controlling the number of unfinished sequences carried over to the next step introduces a tradeoff between efficiency and convergence. A small $\Delta$ reduces overlap and may leave GPUs idle due to tail sequences, while a large $\Delta$ increases overlap but risks inflating per-step latency and introducing staleness in the PPO update.
@@ 4.1 Experimental Setup
We follow state-of-the-art PPO settings using the Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020a). For actor models, we experiment with Qwen2.5-7B, Qwen2.5-7B-Instruct, and Qwen2.5-3B-Instruct,, each augmented with a value head for PPO optimization. The reward model is either a Qwen2.5-7B or a rule-based evaluator (for math tasks). We evaluate on three popular tasks widely used in RLHF research (detailed evaluation setup in Appendix A.1):
  • Free-form generation: Stack-Exchange-Paired (von Werra et al., 2020b), which contains QA pairs with preference labels.
@@
We follow the standard distributed PPO setting. Based on the memory and computation resource requirements of each model, we allocate seven GPUs to the generation and training stages, and one GPU to the scoring stage (i.e., reward model). We compare OPPO against TRL’s PPO (von Werra et al., 2020a), the state-of-the-art and widely adopted framework in PPO. It is worth noting that OPPO is complementary to existing PPO frameworks and can be integrated on top of them. Unless otherwise specified, we use a training batch size of 112.
@@ 4.2 End-to-end Performance Comparison
Figure 3 shows that OPPO consistently accelerates PPO training by $1.8\times$$2.8\times$ across all tasks. On Stack-Exchange with Qwen2.5-7B-Instruct, OPPO reaches a reward of 4.17 in 2,300 minutes versus 4,300 minutes for the baseline, yielding a $1.9\times$ speedup. With Qwen2.5-3B-Instruct on the same dataset, OPPO achieves a reward of 5.12 in 5,200 minutes compared to 13,000 minutes, corresponding to a $2.5\times$ improvement. These gains stem from two sources: (i) intra-step overlap, which hides reward prefilling latency during actor decoding, and (ii) inter-step overlap with dynamic deltas, which mitigates tail stragglers that would otherwise block shorter generations. OPPO achieves $2.4\times$ and $2.8\times$ speedup on OpenCoder-SFT (Stage 2) with Qwen2.5-3B-Instruct, and on GSM8K with Qwen2.5-7B, respectively.
@@
Despite substantial wall-clock speedups, Figure 4 shows that OPPO does not sacrifice training convergence. On Stack-Exchange, OPPO and the baseline follow nearly identical trajectories on both Qwen2.5-7B-Instruct and Qwen2.5-3B-Instruct training, such as reaching a reward of $\sim$2.0 by step 150, then plateauing $\sim$4.1 by step 600 and $\sim$5.12 by step 1,000, respectively. On GSM8K with Qwen2.5-7B, both methods exhibit the same characteristic learning phases: an initial accuracy of 0.70, a dip to 0.66 around steps 25–50 as the model unlearns initial biases, and steady improvement to 0.82 by step 200. Finally, on OpenCoder-SFT (Stage 2) with Qwen2.5-3B-Instruct, both methods converge to a plateau around 2.4 by step 80. Across all tasks, the near-identical step-to-reward curves confirm that OPPO achieves a near-optimal balance between execution efficiency and convergence quality.
@@ 4.3 Ablation Studies
Figure 6 reports their performance on Stack-Exchange-Paired. For Qwen2.5-7B-Instruct, the TRL baseline requires 4,200 minutes to reach a reward of 4.17. Adding only intra-step overlap reduces this to 3,500 minutes ($1.2\times$ speedup), as streaming hides about 17% of scoring latency within generation. However, the gain is bounded by stragglers from the longest sequences in each batch. Applying only inter-step overlap reduces training time further to 2,700 minutes ($1.6\times$ speedup). For Qwen2.5-3B-Instruct, the TRL baseline requires 13,000 minutes to reach a reward of 5.12. Intra-step overlap reduces this to 10,000 minutes ($1.3\times$ speedup), while inter-step overlap achieves 6,300 minutes ($2.06\times$ speedup). Again, all configurations converge to similar final rewards, confirming that intra- and inter-step overlaps address orthogonal bottlenecks while preserving training quality.
@@
Figure 7(b) shows the effect of chunk size on step speed across different model sizes. Small chunks (100 tokens) incur high scheduling and GPU context-switch overhead, reducing throughput despite high overlap. Moderate chunks (500 tokens) strike the best balance, yielding the fastest step speeds by maximizing overlap while avoiding overhead. Large chunks (1000–3000 tokens) reduce overlap opportunities and push execution closer to sequential mode, causing step speeds to rise again. These results highlight that throughput is highly sensitive to chunk size, and the optimal setting depends on model scale and workload.
@@
Table 1: Evaluation results on core tasks (0-shot) and math tasks (5-shot). We report accuracy (%) for TRL trained models and OPPO trained models, along with the absolute change.

### v2

@@ Abstract
Proximal Policy Optimization (PPO)-based reinforcement learning from human feedback (RLHF) is a widely adopted paradigm for aligning large language models (LLMs) with human preferences. However, its training pipeline suffers from substantial inefficiencies due to sequential multi-model dependencies (e.g., reward model depends on actor outputs) and long-tail response lengths, where a few long responses straggle the stage completion. We present OPPO, a novel, lightweight, and model-agnostic PPO-based RLHF framework that improves training efficiency by overlapping pipeline execution. OPPO introduces two novel techniques: (1) Intra-step overlap, which streams upstream model outputs (e.g., actor model) in right-sized chunks, enabling the downstream model (e.g., reward) to begin prefill while the upstream continues decoding; and (2) Inter-step overlap, which adaptively overcommits a few prompts and defers long generations to future steps, mitigating tail latency without discarding partial work. OPPO integrates easily with existing PPO implementations with a lightweight wrapper. Extensive evaluations show that OPPO accelerates PPO-based RLHF training by $1.8\times$$2.8\times$ and improves GPU utilization by $1.4\times$$2.1\times$ without compromising training convergence.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human preferences. Among RLHF methods, Proximal Policy Optimization (PPO) (Schulman et al., 2017) has been the de facto standard due to its training stability and flexibility across diverse reward models and objectives. Following InstructGPT (Ouyang et al., 2022), PPO remains the standard for online alignment in both research and industry. Recent work shows it outperforms offline methods like DPO on reasoning tasks (Xu et al., 2024), and it supports massive-scale training in modern tool chains (Shen et al., 2024). A standard PPO-based RLHF pipeline involves four models: an actor (policy), a critic (value function), a reference policy (for KL regularization), and a reward model trained on human-labeled preferences. Each training step consists of three sequential stages: (1) Generation: the actor generates responses to prompts; (2) Scoring: responses are evaluated by the critic, reference, and reward models; and (3) Training: the actor and critic models are updated using advantage estimates and gradients.
@@
Despite its effectiveness, PPO-based RLHF faces significant training inefficiencies rooted in its multi-model dependencies. Running and coordinating four LLMs imposes substantial resource requirements, and each stage is constrained by its slowest component. For example, the actor model’s generation suffers from severe long-tail latency: a few long responses can delay downstream stages, such as the reward and value models, leading to idle resources and poor training throughput of the pipeline. As LLMs grow larger and context lengths increase, these bottlenecks worsen,  (Grattafiori et al., 2024) making PPO-based RLHF increasingly costly to train (§2.2).
@@
Recent advances tackle PPO-based RLHF inefficiencies from both algorithmic and system perspectives. On the algorithmic side, methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Group-Relative Policy Optimization (GRPO) (Shao et al., 2024) remove components like the value or reward model. However, these approaches often suffer from instability due to sparse rewards, requiring many rollouts to capture intrinsic advantages, and face task-specific reward design challenges (Feng et al., 2024; Fisch et al., 2025; Chen et al., 2024). Asynchronous RLHF (Noukhovitch et al., 2025) reduces pipeline dependencies but introduces staleness, which can harm convergence (§2.2). In contrast, system-level approaches, such as RLHFuse (Zhong et al., 2025), AReal (Fu et al., 2025), and Verl (Sheng et al., 2025), improve throughput via fine-grained parallelism, for example by collocating models to reduce communication overhead or dynamically scaling GPU resources to match workload demands.
@@
In this paper, we explore a complementary opportunity to accelerate PPO-based RLHF: maximizing execution overlap in the training pipeline. We introduce two novel insights: (1) Intra-step Overlap, which streams tokens from the actor to downstream models, enabling generation and scoring stages to overlap the execution without altering the generated response; and (2) Inter-step overlap, which selective overcommits a few prompts per batch and prioritizes faster completions, deferring stragglers to future iterations to hide tail latency without wasting partial generation.
@@
In this paper, we present OPPO, a lightweight PPO-based RLHF training framework that improves training efficiency via pipeline overlap, minimizing idle time without compromising convergence. OPPO novelly addresses the aforementioned challenges:
  • Intra-step Overlap: While the actor generates responses, OPPO streams newly generated tokens to downstream models (e.g., reward model) in adaptive chunks. This enables incremental prefilling and overlaps the generation and scoring stages. Chunk sizes are automatically adjusted online, based on the leftover resources, to balance overlap against resource contention, preserving algorithm correctness and stability (§3.1).
@@
  • Generalized and lightweight: Our evaluations show that OPPO achieves $1.8\times$$2.8\times$ speedup and improves GPU utilization by $1.4\times$$2.1\times$ for PPO-based RLHF with only a lightweight wrapper, and generalizes to other paradigms such as DPO with similar benefits (§4).
@@
Figure 1: (a) In the existing paradigm, the scoring stage begins only after the response is fully generated. In contrast, (b) the OPPO paradigm interleaves scoring with generation without altering the final responses (intra-step overlap), and carries unfinished overcommitted sequences into the next iteration (inter-step overlap). A batch size of 4 and an overcommitment degree of 2 in illustrations.
@@ 2 Background and Motivation
Figure 2: PPO-based RLHF faces (a) varying resource demands across pipeline stages, and (b) heterogeneous rollout lengths, both of which can produce stragglers that prolong step execution. (c) Existing approaches for asynchronous training risk harming convergence.
@@ 2.1 Background: PPO-based RLHF
Figure 1(a) depicts a single step of a standard PPO-based RLHF pipeline. Given a batch of prompts, the actor model generates output sequences. These are then scored by the reward model, producing scalar rewards that reflect alignment with human preferences. A reference model, typically a frozen copy of the base pretrained model, computes a KL divergence penalty that regularizes the update, discouraging the new policy from drifting too far from the original distribution.
@@
where $r_{t}$ is the reward at step $t,V\left(s_{t}\right)$ is the estimated value of state $s_{t},\gamma$ is the discount factor, and $\lambda$ is the generalized advantage estimation (GAE) parameter. In every step, actor model is updated by optimizing the clipped surrogate objective:
@@ 2.2 Training Inefficiency of PPO-based RLHF
Each model in the RLHF pipeline (e.g., actor, reward, and value models) exhibits distinct computational characteristics. Figure 2(a) compares GPU utilization across three GPU types (A40, A100, H200). Response generation in the actor model is memory-intensive due to autoregressive per-token decoding, resulting in low GPU utilization (<40%), whereas scoring and training stages are relatively compute-intensive (e.g., due to long-context prefilling in scoring). This heterogeneous resource utilization highlights how mismatched compute demands across stages create idle GPU time, motivating designs to scavenge unused resources.
@@
Each step involves updating model weights. A natural strategy to improve throughput is to tolerate training staleness, where the reward model evaluates actor outputs from previous steps instead of synchronizing in real time, as in AReal  (Fu et al., 2025). However, as shown in Figure 2(c), asynchronous training (e.g., with staleness 5) can not only slow step-to-reward convergence but also lowers the post-training model quality, emphasizing the need for careful inter-step synchronization in RLHF pipelines.
@@ 3.1 Overlapping Intra-step Training Pipeline
By dividing actor generation into chunks and streaming them to the reward model, OPPO overlaps the actor decoding stage with the reward prefilling stage, hiding latency and reducing execution bubbles. This design naturally benefits setups where models are placed on separate GPUs, but also improves efficiency when models are colocated, due to their mismatched compute demands (Figure 2(a)). To realize intra-step overlap, OPPO partitions actor outputs into right-sized chunks and streams each chunk to the reward model as it is generated. Scoring proceeds progressively within each PPO step: while the actor decodes the $k$-th chunk, the reward model concurrently processes the prefilling of (k1)(k-1)-th chunk. At the end of the step, the reward model completes prefilling for the final chunk and computes the score based on the entire sequence, whose previous chunks have been processed.
Importantly, this streaming does not alter the response generation $y_{i}$, the policy log-probabilities, or the critic/value terms used in computing the advantage $\hat{A}(y_{i})$. Formally, letting $y_{i}$ denote the full response and $y_{i}^{(1)},\dots,y_{i}^{(T_{i})}$ its prefixes with $y_{i}^{(T_{i})}=y_{i}$, the streamed gradient estimator is
@@ 3.2 Overlapping Inter-step Training Pipeline
While intra-step overlap improves efficiency within a single PPO step, it does not fully address tail latency caused by the heterogeneous response lengths within a batch. Here, the response must complete generation before reward scoring and subsequent policy updates. Due to the long-tailed distribution of generation lengths, a few slow prompts can delay the entire step. This motivates an inter-step design that allows overlapping across PPO steps without hurting convergence.
@@
The overall procedure, combining intra- and inter-step overlap, is summarized in Algorithm 1, where the buffer holds up to $B+\Delta$ sequences, and generation proceeds in parallel with intra-step streaming. The threshold $\Delta$, which controlling the number of unfinished sequences carried over to the next step, introduces a tradeoff between efficiency and convergence. A small $\Delta$ reduces overlap and may leave GPUs idle due to tail sequences, while a large $\Delta$ increases overlap but risks inflating per-step latency and introducing staleness in the PPO update.
@@ 4.1 Experimental Setup
We follow state-of-the-art PPO settings using the Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020). For actor models, we experiment with Qwen2.5-7B, Qwen2.5-7B-Instruct, and Qwen2.5-3B-Instruct, each augmented with a value head for PPO optimization. The reward model is either a Qwen2.5-7B or a rule-based evaluator (for math tasks). We evaluate on three popular tasks widely used in RLHF research (detailed evaluation setup in Appendix A.1):
  • Free-form generation: Stack-Exchange-Paired (von Werra and others, 2020), which contains QA pairs with preference labels.
@@
We follow the standard distributed PPO setting. Based on the memory and computation resource requirements of each model, we allocate seven GPUs to the generation and training stages, and one GPU to the scoring stage (i.e., reward model). We compare OPPO against TRL’s PPO (von Werra et al., 2020), the state-of-the-art and widely adopted framework in PPO. It is worth noting that OPPO is complementary to existing PPO frameworks and can be integrated into them. Unless otherwise specified, we use a training batch size of 112.
@@ 4.2 End-to-end Performance Comparison
Figure 3 shows that OPPO consistently accelerates PPO training by $1.8\times$$2.8\times$ across all tasks. On Stack-Exchange with Qwen2.5-7B-Instruct, OPPO reaches a reward of 4.17 in 2,300 minutes versus 4,300 minutes for the baseline, yielding a $1.9\times$ speedup. With Qwen2.5-3B-Instruct on the same dataset, OPPO achieves a reward of 5.12 in 5,200 minutes compared to 13,000 minutes, corresponding to a $2.5\times$ improvement. These gains stem from two sources: (i) intra-step overlap, which hides reward prefilling latency during actor decoding, and (ii) inter-step overlap with dynamic $\Delta$, which mitigates tail stragglers that would otherwise block shorter generations. OPPO achieves $2.4\times$ and $2.8\times$ speedup on OpenCoder-SFT (Stage 2) with Qwen2.5-3B-Instruct, and on GSM8K with Qwen2.5-7B, respectively.
@@
Despite substantial wall-clock speedups, Figure 4 shows that OPPO does not sacrifice training convergence. On Stack-Exchange, OPPO and the baseline follow nearly identical trajectories on both Qwen2.5-7B-Instruct and Qwen2.5-3B-Instruct training, such as reaching a reward of $\sim$2.0 by step 150, then plateauing at $\sim$4.1 by step 600 and $\sim$5.12 by step 1,000, respectively. On GSM8K with Qwen2.5-7B, both methods exhibit the same characteristic learning phases: an initial accuracy of 0.70, a dip to 0.66 around steps 25–50 as the model unlearns initial biases, and steady improvement to 0.82 by step 200. Finally, on OpenCoder-SFT (Stage 2) with Qwen2.5-3B-Instruct, both methods converge to a plateau around 2.4 by step 80. Across all tasks, the near-identical step-to-reward curves confirm that OPPO achieves a near-optimal balance between execution efficiency and convergence quality.
@@
TRL OPPO
Mean latency (s) 498.30 111.08
Speed up 1.00x 4.49x
[…]
OPPO improves performance in multi-node settings.
[…]
OPPO delivers improvements over different model parallelism plans.
[…]
@@ 4.3 Ablation Studies
Figure 6 reports their performance on Stack-Exchange-Paired. For Qwen2.5-7B-Instruct, the TRL baseline requires 4,200 minutes to reach a reward of 4.17. Adding only intra-step overlap reduces this to 3,500 minutes ($1.2\times$ speedup), as streaming hides about 17% of scoring latency during generation. However, the gain is bounded by stragglers from the longest sequences in each batch. Applying only inter-step overlap reduces training time further to 2,700 minutes ($1.6\times$ speedup). For Qwen2.5-3B-Instruct, the TRL baseline requires 13,000 minutes to reach a reward of 5.12. Intra-step overlap reduces this to 10,000 minutes ($1.3\times$ speedup), while inter-step overlap achieves 6,300 minutes ($2.06\times$ speedup). Again, all configurations converge to similar final rewards, confirming that intra- and inter-step overlaps address orthogonal bottlenecks while preserving training quality.
Deferred steps 0 1 2 3 Avg. deferred steps
Share of requests 78.48% 20.20% 0.23% 1.05% 0.24
Table 2: Distribution of requests deferral shows most requests are not deferred, and nearly all others are delayed by only a single step.
Robustness and Staleness.
As detailed in Algorithm 1, the $\Delta$ controller adapts to a windowed reward trend, updating $\Delta$ only through bounded, gradual steps. This design prevents abrupt jumps and effectively filters short-term oscillations. The request-deferral distribution in Table 2 confirms the stability of this approach: the vast majority of requests are processed immediately, and nearly all deferred requests are delayed by only a single step. This indicates neither perpetual deferral of difficult prompts nor excessive staleness that would affect rewards.
@@
Figure 7(b) shows the effect of chunk size on step speed across different model sizes. Small chunks (100 tokens) incur high scheduling and GPU context-switch overhead, reducing throughput despite high overlap. Moderate chunks (500 tokens) strike the best balance, yielding the fastest step speeds by maximizing overlap while avoiding overhead. Large chunks (1000–3000 tokens) reduce overlap opportunities and push execution closer to sequential mode, causing step times to rise again. These results highlight that throughput is highly sensitive to chunk size, and the optimal setting depends on model scale and workload.
@@
Table 3: Evaluation results on core tasks (0-shot) and math tasks (5-shot). We report accuracy (%) for TRL-trained models and OPPO-trained models, along with the absolute change.
Applicability beyond PPO.
OPPO’s benefits extend to any online preference-optimization method involving variable-length on-policy generations (e.g., DPO or GRPO). These methods can adopt the same scheduling logic: generate $B{+}\Delta$ items, update on the first $B$ completions, and carry unfinished long generations forward to the next iteration. This strategy reduces tail latency without altering the optimization objective or the distribution of responses used for updates. As shown in Figure 3, OPPO achieves $2.35\times$ faster convergence in a rule-based PPO setting on GSM8K (without a reward model), confirming that the inter-step overlap mechanism remains effective even in non-standard or simplified RLHF pipelines.
VeRL w/ DP VeRL w/ DP+SP AReaL OPPO
Mean latency (s) 125.36 120.47 109.92 99.84
Table 4: OPPO achieves the lowest per-step latency under identical hardware and rollout settings, suggesting system-level benefits beyond VeRL (DP, DP+SP, Fully Async w/ SP) and AReaL.